OpenStudy (anonymous):

Consider the line that passes through the point (3, -2) and has a slope of 2. Part 1: Write the equation of this line using point-slope form. (2 points) Part 2: Using your equation from part 1, rewrite this equation in slope-intercept form. Make sure to show all of your work. (2 points)Part 3: Using your equation from part 2, rewrite this equation in standard form. Make sure to show all of your work. (2 points)

OpenStudy (anonymous):

Not Asking for the answer just like a step-through please? New to this stuff and confusing when school just throw's it at you =/.

OpenStudy (anonymous):

Point slope formula is y-y=ms(x-x)

OpenStudy (anonymous):

Wait ill work it out.

OpenStudy (anonymous):

It's so confusing to me =/ I don't know what to do.

OpenStudy (anonymous):

The first part is y-(-2)=2(x-3) when you substitute 2 to x and three you will get y-(-2) = 2x-3 and then you subtract -2 from both sides to get y=2x-5 ànd that's in slope intercept form and then

OpenStudy (anonymous):

I can do the next part =3 THANK YOU! just need help on the first part

OpenStudy (anonymous):

How do I reward you a medal =3 new to this =D but I really appreciate you taking time out of you're life to help me =3.

OpenStudy (anonymous):

To put it in standard form subtract 2x from y to get -2x+y=5

OpenStudy (anonymous):

Oh okay and I don't know lol

OpenStudy (anonymous):

First part was like whoa a shocker =D first time ever seen